OpenStudy (anonymous):

while shopping for clothes, consuelo spent $3 less than twice what brenda spent. Consuelo spent $59. how much did Brenda spend?

OpenStudy (anonymous):

$26.50

OpenStudy (anonymous):

let c = consuelo's spent money and b = brenda's spent money c = 2b-3 = 59 or simply 2b-3 = 59 solve for b

OpenStudy (anonymous):

so add 3 to both sides?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

$31

OpenStudy (anonymous):

Lol Woah I Might Need Help Then Too Lol...

OpenStudy (anonymous):

That's what I had :) IDK about the other answer

OpenStudy (anonymous):

I simply divided 59 by 2 and subtracted 3 from the total.

OpenStudy (anonymous):

2B - 3 = 59 -> B = ( 59 + 3 )/2 = $31 ( Brenda spent)

OpenStudy (anonymous):

oops i meant 62/2=31